OpenStudy (anonymous):

Find f(-2) if f(x)=x⁴+2x²-1

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = x^4+2x^2-1\] \[\Large f(-2) = (-2)^4+2(-2)^2-1\] \[\Large f(-2) = 16+2(4)-1\] \[\Large f(-2) = 16+8-1\] \[\Large f(-2) = 23\]